OpenStudy (anonymous):

Twenty persons are invited for a party. In how many different ways can the host be seated at a circular table, if two particular persons are to be seated on either side of the host?

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

two persons so 2!

OpenStudy (paki):

do you know permutation....?

OpenStudy (paki):

@dg2

OpenStudy (anonymous):

somewhat!blindly without proper concept

OpenStudy (anonymous):

@paki i hav lost the connection

OpenStudy (anonymous):

19 ways twenty persons seated,then two persons on either side so 2!

OpenStudy (anonymous):

18!*2 ways i know the answer but it is not clear to me

OpenStudy (paki):

what is the total number of people...?

OpenStudy (anonymous):

20

OpenStudy (queelius):

I don't understand where there are only 18!*2 ways either. Are you sure you didn't mean 18!*20?

OpenStudy (queelius):

where=why

OpenStudy (anonymous):

yes

OpenStudy (queelius):

Wait, just realized I made a mistake. Let me reconsider.

OpenStudy (queelius):

First, is the host included among those 20 people invited?

OpenStudy (anonymous):

20 persons so 19 ways

OpenStudy (queelius):

Okay, assuming the host is one of the 20 people. Here's my formulation. First, of the 20 seats, choose 1 for the host. Next, he has two seats next to him. For the 2 people that are suppose to sit next to him, there are 2 ways to seat them. Now, we have 17 remaining people, and they can be in any arrangement in the remaining chairs. 20*2*17!

OpenStudy (anonymous):

After fixing the places of three persons (1 host + 2 persons) and treating them as 1 unit we can arrange the total (20 - 2 + 1) = 19 units in 18! ways. Again these particular persons can sit on either side of the host in 2 ways. Hence the total number of ways is 18! × 2

OpenStudy (queelius):

First, you are not counting the host as 1 of the 20?

OpenStudy (anonymous):

then

OpenStudy (queelius):

Then what? I'm just trying to figure out if you're including the host as one of the twenty. :)

OpenStudy (anonymous):

see 20 people seated in how many ways

OpenStudy (queelius):

Okay, since I assume this means we are to count the host as one of the twenty people, my earlier formation would seem correct: 20 ways to seat the host, 2 ways to seat the people who are suppose to sit next to him, and 17! ways to seat the rest. So, 20*2*17! = 30*17!

OpenStudy (anonymous):

nope .18!*2 is the answer

OpenStudy (queelius):

How do you think they arrived at 18!*2?

OpenStudy (queelius):

First, it seems I was solving the wrong problem. I was solving the problem: How many ways can people be seated at the circular table, given that the host will have two particular people of those 19 non-hosts who must sit next to him.

OpenStudy (anonymous):

After fixing the places of three persons (1 host + 2 persons) and treating them as 1 unit we can arrange the total (20 - 2 + 1) = 19 units in 18! ways. Again these particular persons can sit on either side of the host in 2 ways. Hence the total number of ways is 18! × 2

OpenStudy (queelius):

Oh well, I don't think I understand what the question is actually asking. The answer makes no sense to me.

OpenStudy (anonymous):